The sum of 2 two-digit numbers is 89 and their difference is 19. What are the two numbers?






Answer by Cromlix(4381)

x + y = 89
x - y = 19
Add
2x = 108
x = 54
Substitute x = 54
into:
x + y = 89
54 + y = 89
y = 89 - 54
y = 35
54 and 35
Hope this helps.
:-)